### v1

# Mirror-enhanced acceleration induced geometric phase: towards detection of Unruh effect

Abstract
[…]
Introduction.– The Fulling-Davies-Unruh effect [1, 2, 3, 4] plays a crucial role in understanding the quantum fields in curved spacetime. A uniformly accelerated observer sees a thermal bath in the vacuum of an inertial observer. It also closely relates to the so-called Hawking radiation [5, 6] through the equivalence principle. The accelerated observer is analogous to the static observer far away from the black hole, and the static observer is analogous to the free-falling observer near the black hole horizon. In this way, detecting the Unruh effect backs the existence of the Hawking effect. Hence it may shed light on the quantum nature of strong gravity. However those phenomenon are far from experimental verification. Despite its importance, requirement of significantly high acceleration ($\sim 10^{21}m/s^{2}$ for $1$ Kelvin temperature [4]) hindering its experimental detection. Numerous efforts have been made to detect the Unruh effect in several approaches [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. However either the proposed linear acceleration is too high ($\sim 10^{9}m/s^{2}$ with qubit energy gap $\sim$ MHz [13]) or investigation has been done with circular motion (see e.g. [20] where required acceleration is reported as $\sim 10^{7}m/s^{2}$ with qubit energy gap $\sim 10$ MHz, rotating frequency $\sim 0.1$ MHz and radius of path $\sim 10^{-3}m$).
The adiabatic-cyclic evolution may generate a geometric phase (known as Pancharatnam-Berry phase (PBP)) in a quantum system [21, 22]. It contains information of systems’ motion and background environments [23, 24, 25, 26]. Also, interferometric set-ups require significantly less acceleration to detect the Unruh phenomenon using the phase [23, 24]. In this set-up, one employs the two Unruh-DeWitt (UD) detectors: one detector accelerates in one arm of the interferometer, and another inertially moves in the other arm. Both of the detectors accumulate some PBP depending on their motions. Detection of this phase difference can reveal an indirect observation of the Unruh effect. In free space, the phase difference is experimentally realisable with minimum acceleration $\sim 10^{17}m/s^{2}$ [23, 24].
In a recent study, we observed that the detector’s transition rate can get reasonably modified in presence reflecting boundaries (called as mirrors) [27]. PBP is closely related to the detector’s transition rate [23, 28, 20], and depends on the boundary [29]. Moreover, the phase is much more sensitive than the transition rate from an experimental point of view. Therefore, PBP may be susceptible to low accelerations in the presence of mirrors.
Motivated by these facts here we aim to study PBP difference (denote as $\delta\phi_{B}$) between an linearly accelerated UD detector and the inertial one in presence of mirrors. The acceleration is along one of the directions parallel to mirror’s plane. We investigate two different situations: the detectors are in presence of (i) single mirror and (ii) in between double parallel mirrors, respectively. The first situation shows that when the detectors are at the same distance from the mirror, $\delta\phi_{B}$ is at the measurable range ($\gtrsim 5.27\times 10^{-6}\,rad$ [30]) for much lower accelerations depending on the distance ($z_{0}$) from the mirror and the atom’s energy gap. For the analysis, one may utilize tuneable flux qubits as detectors. With the energy gap of the qubit in GHz order and $z_{0}\sim 10^{1}-10^{4}\,m$, we observe a measurable phase difference with acceleration as low as $a\sim 10^{10}m/s^{2}$. For the double mirrors, we do similar set-ups. The atom is allowed to move within the space in between the mirrors. Here $\delta\phi_{B}$ depends on $z_{0}$, separation distance between the two mirrors $L$ and the atom’s energy gap. We can achieve measurable phase differences for acceleration as low as $a\sim 10^{9}m/s^{2}$ with suitable choices of parameters (e.g. $z_{0}\sim 10^{1}\,m$, $L\sim 10^{2}\,m$ and energy gap $\sim 1$ GHz). However qubits with energy gap in MHz order (as considered in [24, 13, 20]), we find the required acceleration for two mirrors brings down by considerable amount $a\sim 10^{6}m/s^{2}$. This is not only much less than the earlier proposals [24, 13, 20], but also is very much achievable within the current technology (e.g. using thermal gradient technique [31]).
Let us now proceed towards the calculation to justify our claim.
The UD detector model and PBP.– Consider the total Hamiltonian of the system $\hat{H}=\hat{H}_{d}+\hat{H}_{\phi}+\hat{H}_{\text{int}}$. Here $\hat{H}_{d}$ is the Hamiltonian of the detector, is taken to be $\hat{H}_{d}=\frac{\hbar}{2}\omega_{0}\hat{\sigma}_{3}$, where $\hat{\sigma}_{3}$ is Pauli spin matrix and $\hbar\omega_{0}$ is the energy gap of the detector. $\hat{H}_{\phi}$ is the background field’s (real scalar field $\hat{\phi}$) Hamiltonian and $\hat{H}_{\text{int}}$ is the interaction Hamiltonian, given by $\hat{H}_{\text{int}}=\lambda\,\hat{m}\,\hat{\phi}$. Here, $\hat{m}$ is the operator (chosen to be as $\hat{m}=(1/2)\sum_{i}\hat{\sigma}_{i}$) representing the detector, which is analogous to the dipole operator corresponding to electro-magnetic interactions and $\lambda$ is the coupling constant. Initially, the field is considered to be in the Minkowski vacuum state, and the detector’s state is $\hat{\rho}(0)=|\psi(0)\rangle\langle\psi(0)|$. With perturbative analysis, the reduced density matrix at a later time can be found by solving the Kossakowski-Lindblad equation [32, 33, 34]
$\frac{\partial\hat{\rho}(\tau)}{\partial\tau}=-\frac{i}{\hbar}[\hat{H}_{\text{eff}},\hat{\rho}(\tau)]+\hat{\mathcal{L}}[\hat{\rho}(\tau)]\,,$ (1)
where, $\hat{H}_{\text{eff}}=\frac{\hbar}{2}\Omega\hat{\sigma}_{3}$ is the effective Hamiltonian with well-known Lamb shift correction ($\Omega=\omega_{0}+\omega_{L}$$\omega_{L}$ is the Lamb shift frequency), and
^$\hat{\mathcal{L}}[\hat{\rho}]=\frac{1}{2}\sum_{i,j=1}^{3}a_{ij}[2\hat{\sigma}_{j}\hat{\rho}\hat{\sigma}_{i}-\hat{\sigma}_{i}\hat{\sigma}_{j}\hat{\rho}-\hat{\rho}\hat{\sigma}_{i}\hat{\sigma}_{j}]\,,$ (2)
where $a_{ij}=A(\delta_{ij}-\delta_{i3}\delta_{j3})-iB\epsilon_{ijk}\delta_{k3}$, and $\tau$ is the detector’s proper time.
With consideration of the initial detector state $|\psi(0)\rangle=\cos\frac{\theta}{2}|+\rangle+\sin\frac{\theta}{2}|-\rangle$, we obtain the time-dependent reduced density matrix (by tracing out the field’s degrees of freedom) as [23]
$\hat{\rho}(\tau)=\begin{pmatrix}f(\tau)&\frac{1}{2}e^{-2A\tau-i\Omega\tau}\sin\theta\\ \frac{1}{2}e^{-2A\tau+i\Omega\tau}\sin\theta&1-f(\tau)\end{pmatrix}~{},$ (3)
where $f(\tau)=e^{-4A\tau}\cos^{2}\frac{\theta}{2}+\frac{B-A}{2A}(e^{-4A\tau}-1)$, $A=\frac{1}{4}[\gamma(\omega_{0})+\gamma(-\omega_{0})]$ and $B=\frac{1}{4}[\gamma(\omega_{0})-\gamma(-\omega_{0})]$. The functions $\gamma$’s are basically detectors emission or excitation rates, given by
$\gamma(\omega_{0})=\frac{\lambda^{2}|\langle+|\hat{m}|-\rangle|^{2}}{\hbar^{2}}\int_{-\infty}^{\infty}d(\Delta\tau)\,e^{i\omega_{0}\Delta\tau}\,G_{W}(\Delta\tau)\,,$ (4)
where $G_{W}(\Delta\tau)$ is positive frequency Wightman function for $\hat{\phi}$. By following [23], one can solve this density matrix for the eigenstates and the PBP turns out to be
$\phi_{B}=-\pi(1-\cos\theta)-\frac{2\pi^{2}B\sin^{2}\theta}{\omega_{0}}\left(2+\frac{A}{B}\cos\theta\right)~{}.$ (5)
Evaluation of $A,B$ will provide $\phi_{B}$.
Here we consider the detector is accelerating in $x$-direction whose trajectory is given by
$t=\frac{c}{a}\sinh\frac{a\tau}{c},~{}x=\frac{c^{2}}{a}\cosh\frac{a\tau}{c},~{}y=0~{},z=z_{0}~{},$ (6)
and the two mirrors are at $z=0$ and $z=L$ (with $0<z_{0}<L$). $G_{W}(\Delta\tau)$ in presence of two mirrors (infinitely extended) is given by [35, 36, 27]
$G_{W}(\Delta\tau)=-\frac{\hbar}{4\pi^{2}c}\sum_{n=-\infty}^{\infty}\left(\frac{1}{(c\Delta{t}-i\epsilon)^{2}-\Delta{x}^{2}-(2L\,n)^{2}}\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left.\frac{1}{(c\Delta{t}-i\epsilon)^{2}-\Delta{x}^{2}-(2z_{0}-2L\,n)^{2}}\right)\,.$ (7)
By construction, this green function has symmetry between two points $(z_{0},~{}L-z_{0})$. Note that consideration of only $n=0$ term of the above leads to the same for the single mirror situation, located at $z=0$. So in future whenever discussion on the single mirror case will be done, we will take only the $n=0$ term from the expressions, evaluated for double mirrors.
For our case, we find
$A=\frac{\kappa a}{16\pi c}\coth\left(\frac{\pi\omega_{0}c}{a}\right)\sum_{n=-\infty}^{\infty}\left(J(Ln)-J(Ln+z_{0})\right)~{};$
$B=\frac{A}{\coth\left(\frac{\pi\omega_{0}c}{a}\right)}~{},$ (8)
where $J(u)=\frac{\sin\left(\frac{2\omega_{0}c}{a}\sinh^{-1}(|\frac{au}{c^{2}}|)\right)}{|\frac{au}{c^{2}}|\sqrt{\left(\frac{au}{c^{2}}\right)^{2}+1}}$ and $\kappa=\frac{\lambda^{2}|\langle+|\hat{m}|-\rangle|^{2}}{\hbar c^{3}}$. Considering only $n=0$ term, one gets $A$ and $B$ in presence of single mirror. Discarding the $z_{0}$-dependent term from these latter $A$ and $B$, we obtain those for free space. For the double-mirror system the summation over $n$ goes from $\infty$- ∞ to $\infty$. However for very large values of $n$, (when $a(Ln+z_{0})\approx aLn$), two terms inside the summation will cancel each other ($J(Ln)-J(Ln+z_{0})\approx 0$). Therefore for sufficiently large values of $n$, we may set a cut off on upper and lower limit of $n$ in order to evaluate the summation in (Mirror-enhanced acceleration induced geometric phase: towards detection of Unruh effect). Hence in numerical calculation we choose max$|n|$ to a large finite value. In this set-up the possibility to measure minimum PBP ($\sim 10^{-6}~{}rad$) with low acceleration ($a$) can be realized as follows. For very low value of $a$ we have $\coth\frac{\pi\omega_{0}c}{a}\approx 1$ and hence (Mirror-enhanced acceleration induced geometric phase: towards detection of Unruh effect) implies $A\approx B$. Therefore in this situation the relations for $A$ and $B$ with emission rate $\gamma(\omega_{0})$ and absorption rate $\gamma(-\omega_{0})$ of the atom implies $\gamma(-\omega_{0})\approx 0$. Then $\delta\phi_{B}$ (see (5)) mostly depends on $\gamma(\omega_{0})$. Moreover $A(\approx B)$ is mostly dominated by the term contributed by the effects of mirror(s) (see the term with summation in (Mirror-enhanced acceleration induced geometric phase: towards detection of Unruh effect)). Therefore controlling suitable values of the parameters, like $z_{0}$, $L$, it may be possible to reach at the desired goal.
Parameters towards detecting PBP.– We need to define dimensionless parameters for numerical analysis of the geometric phase. Denote dimensionless acceleration and distances as $a/\omega_{0}c$, $L\omega_{0}/c$ and $z_{0}\omega_{0}/c$, respectively. Therefore, one can see the quantities $\coth\left(\frac{\pi\omega_{0}c}{a}\right)$ and $J(Ln)$ (or $J(Ln+z_{0})$) are dimensionless. The constant $\kappa$ is also dimensionless, since $\lambda\hat{m}\hat{\phi}$ has to be dimensionally same as $\hbar\omega$ ($\hat{\phi}$ has dimension of $\sqrt{\hbar/\omega L^{3}}$). Thus, as expected phase $\phi_{B}$ is dimensionless.
Our analysis aims to evaluate the effect of acceleration in spaces with or without mirrors. Therefore the object of our interest is $\delta\phi_{B}=\phi_{B}-\lim_{a\to 0}\phi_{B}$ for double, single and no-mirror systems. Also note that, in SI unit $\hat{H}_{d}\sim\hbar\omega_{0}\sim 10^{-25}$ with $\omega_{0}\sim 1$ GHz. In perturbative calculations, one must have $\hat{H}_{int}<\hat{H}_{d}$, or $\lambda\hat{m}\hat{\phi}<10^{-25}$. Alternatively, we can also write $\lambda^{2}|\langle+|\hat{m}|-\rangle|^{2}\langle\hat{\phi}\hat{\phi}\rangle<10^{-50}$. We know $G_{W}\sim\hbar/c\sim 10^{-42}$, thus $\lambda^{2}|\langle+|\hat{m}|-\rangle|^{2}<10^{-8}$ and hence $\kappa<10^{2}$. Therefore for the numerical analysis we choose $\kappa\approx 1$.
Figure 1: We plotted $\delta\phi_{B}$ with respect to $\omega_{0}c/a$, for $\omega_{0}L/c=10$ and $\theta=\pi/4$. Solid, dashed and dotted lines are for double, single and no-mirror (independent of $\omega_{0}z$) systems, respectively.
In a currently achievable experimental set-up, one can measure a phase difference with precision as small as $5.27\times 10^{-6}\,rad$ [30]. Also, one can implement tunable flux qubits, which can have energy gap in the range 111110101010 GHz [37, 38, 39, 38, 40]. For our purpose, we use $\theta=\pi/4$ and $\omega_{0}=1$ GHz. With this choice, the dimensionless acceleration $a/\omega_{0}c=1$ implies $a=3\times 10^{17}m/s^{2}$, and $\omega_{0}z_{0}/c=1$ implies the distance $z_{0}=0.3m$. In Fig. 1, we have shown $\delta\phi_{B}$ double, single and no-mirror systems in solid, dashed and dotted lines, respectively. Free space $\delta\phi_{B}$ decays sharply around $a/\omega_{0}c=1$ (i.e., $a=3\times 10^{17}m/s^{2}$). At $a/\omega_{0}c=0.1$, $\delta\phi_{B}$ is in order $\sim 10^{-16}\,rad$; exceptionally below the experimentally observable range. However, in that acceleration, $\delta\phi_{B}$ for the double and single mirror systems are in the observable range. This indicates the possibility of observing the phase difference in much lower accelerations.
Figure 2: We plotted $\delta\phi_{B}$ for single-mirror system with respect to $\omega_{0}z_{0}/c$ for (a) $a/\omega_{0}c=10^{-2}$, (b) $a/\omega_{0}c=10^{-3}$ and (c) $a/\omega_{0}c=10^{-7}$ (range of $\omega_{0}z_{0}/c$ is chosen where $\delta\phi_{B}$ has maximum values), respectively with the choice $\theta=\pi/4$.
[…]
Figure 3: We plotted $\delta\phi_{B}$ for double-mirror system with respect to $\omega_{0}z_{0}/c$ for (a) $L\omega_{0}/c=10.0$ and (b) $L\omega_{0}/c=500.0$ and $1000.0$, respectively with $\theta=\pi/4$.
We next explore the same in presence of a double mirrors. In Fig. 3, we have plotted $\delta\phi_{B}$ with respect to $\omega_{0}z_{0}/c$ for $L\omega_{0}/c=10.0$ ($L=3m$), $500.0$ ($L=150m$) and $1000.0$ ($L=300m$), respectively. We choose $\omega_{0}=1$ GHz and $\theta=\pi/4$. The first subfigure with $L\omega_{0}/c=10.0$ shows that $\delta\phi_{B}$ is in the experimentally observable range when accelerations $a/\omega_{0}c\sim 10^{-5},~{}10^{-6}$ and $10^{-7}$ (i.e. $a\sim 10^{12},~{}10^{11}$ and $10^{10}m/s^{2}$, respectively). Second subfigure with $L\omega_{0}/c=500.0$ and $1000.0$ shows the similar nature with $a/\omega_{0}c\sim 10^{-8}$ ($a\sim 10^{9}m/s^{2}$). Both subfigures show that the highest peak of $\delta\phi_{B}$ decreases as we use lower values of accelerations. Note that for tiny accelerations, the value $\max|n|$ of the summation has to be much larger to satisfy $J(Ln)\approx J(Ln+z_{0})$. To find out the appropriate $\max|n|$ value, we checked the values of $\delta\phi_{B}$ for $\max|n|$ from $0.1$ to $1\,million$ (in intervals of $0.1\,million$) and used the particular $\max|n|$ value, after which $\delta\phi_{B}$ approximately remain constant. For our parameter values $a/\omega_{0}c\sim 10^{-5},~{}10^{-6},~{}10^{-7}$ and $10^{-8}$, we used value of $\max|n|=0.1,~{}0.4,~{}0.4$ and $0.5\,million$, respectively. In all the above cases, we used the energy gap of the qubit to be $1$ GHz. However for $\omega_{0}$ in MHz order (as mentioned in [24, 13, 20]), the required acceleration can be brought down up to $10^{6}m/s^{2}$ with $\omega_{0}=1$ MHz, $L=150$ (or $300\,km$), $z_{0}=3.3\,km$. This required acceleration is significantly lower than other configured systems in the literature [13].
Conclusion.– Within this simple model, we observed that the Unruh effect can be detected with much lower accelerations using the PBP in the presence of reflecting boundaries. With a single mirror, the effect can be observed with acceleration as low as $\sim 10^{10}m/s^{2}$, with a significant separation between the atom and the boundary. With shorter separations, one can still detect the phase with accelerations of a few orders less than that of a free space. With double-mirrors, the effect is more apparent and situation is much more inspiring to build experimental setup. Here, one needs acceleration to detect the phase can be as low as $\sim 10^{9}m/s^{2}$ depending on the separation between the boundaries and the atom’s position inside them for energy gap $\sim$ GHz. The order of the required acceleration can be $10^{3}$ times less if one can utilize a qubit with the energy gap in MHz order. Also, further reducing the acceleration is possible by exploring other possible parameter ranges with enhanced numerical capabilities.
Our study shows that boundaries can significantly influence the geometric phase with much lower accelerations. In fact the minimum observable phase can be detected with very low acceleration when two mirrors are kept within a finite distance. The required acceleration is very low compared to the earlier proposals and hence can be a very strong candidate for construction of an apparatus to detect Unruh effect. However, looking at the strong effects of the mirrors, we may expect cavity arrangement may show further improvement in the PBP. Moreover here we did not incorporate the effects of edges of the mirrors. Incorporation of all these may be very important in this context. We keep these investigation for future.
Acknowledgements.
Acknowledgments.– DB would like to acknowledge Ministry of Education, Government of India for providing financial support for his research via the PMRF May 2021 scheme. The research of BRM is partially supported by a START-UP RESEARCH GRANT (No. SG/PHY/P/BRM/01) from the Indian Institute of Technology Guwahati, India.